Term Goals Fall 2015 for Research Support and Special Projects

General environment

File storage

  • Meet with Dave and Anthony to review plans for Distributed File Service - Lawrence/Lori w/Dave/Anthony - November 2015
    • Milestone: meet and make progess
    • Met Oct. 28th - made notes in ST#98307
    • Lori has updated quotes
    • Dave/Lawrence/Lori to create updated proposal for January 2016

Networking

Firewall

  • Finish group consultations - Early December 2015 - Lawrence / RSG staff
    • Overall: 88134
      • -> some progress, but still a number of groups to confirm upcoming changes
    • CHIL: 99192 - Gordon
    • NPSG: 99194 - Gordon
      • met with Gladimir - 2015-11-05
    • Lori: affected systems: Syn, Daytona - currently on 26, which will only have VPN,SSH and web by default
    • Lori: Ripple - 129.97.119.223/26 (VLAN 500) - needs SSH (so ssh and web would be fine) - 102465
  • Move CrySP grads to VLAN 7/26 - Gordon

CSCF Special Projects

CSCF internal services

Machine Rooms

DC 3556 (Research Machine Room)

  • Finish Insurance Claim - November 2015 - Lawrence
    • 95303
      • submitted to insurance, awaiting final approval (and cheque!)

DC 3558 (CS Infrastructure Room)

  • Move all RSG equipment out of Row B- December 2015 - Lawrence/RSG
  • 101485
    • all done as of 2015-12-11

Monitoring

  • Report when server hits predefined temperature limit - Gordon
    • 97155
    • Note: confirm with PLG that they're ok with them paying for it (confirmed - see ST#97155, original email from Peter asking about these sensors)
    • -> no progress

Research computing

Graduate student workstations

  • Streamline the AD join process - Mike / Clayton
    • 101333
    • Milestone: revised script that works with current image for Winter 2016 post-install steps
    • Future: (Winter 2016) Update for Spring/Fall 2016 image
      • Mike says that he has done some work in improving the AD script. Need to document in above ST or note related ST

HPC

generic cluster ("paper")

  • Rack hardware we have in DC 3558 A1 - Lawrence/Lori/Mike/CSCF Coop
    • 99112
    • Report time spent on this project
    • -> awaiting rack to be emptied

Ganglia Portal

  • ST#98196
    • waiting on VM to be deployed (Dave - can you make that happen?)

Visitor loaner equipment / Researcher-owned machines

  • Workstation imaging: Streamline via Clonezilla - Click'N'Go - Mike
    • ST#102946
    • involves Dave and Phil (See Also ST#97089 for knowledge transfer)
    • -> Mike has done a lot of work on this process and it seems to be ready, need to ensure that this is shared with Phil and integrated with their procedures
  • Milestone: unclear

Administration

Billing

  • Fall 2015 bills by end of November
    • not yet done, needs to happen very soon
  • Future: Dave would like Winter bills done by late January/early February

Research Groups

AI

  • Setup guest computers - Mike
    • Mike has worked with John (our co-op) to review existing machines and update/replace as needed

BIF

  • rebuild m160 File storage - Fall 2015 - Mike / Lori
    • Update (2015-11-05)
      • Mike has discussed with students and Ming - and they are excited to have this happen
      • need to transfer all of the data over to Asimov (or other storage server)
  • memory upgrade for all of the nodes - planning to be done in November (ST#99218)
    • -> done

Boutaba

  • Management of CN cluster - Ronaldo
  • Documentation of CN cluster - Ronaldo
    • the above are waiting for the return of one of KEshav's grad students away Fall 2016, returning in January

Brecht

  • Migrate rocket to new server - Ronaldo / Lori
    • Lori - ST#101572 - in progress

Cabernet

  • Upgrade cluster OS to 14.04 (priority)
  • Upgrade GPUs? Lori/Mike

Daytona

  • Reorganize, add 2 nodes - December 2015 - Lori
  • 100216
    • in progress, should be complete early January. Some systems blocking the expansion are being moved the week of Dec 15th

Games Institute

  • Fileserver / Backup server installation - November 2015 - Lori
  • 97715
    • done

HCI

  • Rebuild NAS - Ronaldo
    • done

Himrod

  • Monitoring/Reporting/OS imaging - Lori

ISS4E

  • Move backup system to new hardware - Ronaldo
  • Management of the servers - Ronaldo
  • Documentation - Ronaldo
  • Setup raindrop4.cs in ERC - Ronaldo
    • [[https://cs.uwaterloo.ca/cscf/internal/request_debug/UpdateRequest?101439][101439
  • -> done

PLG

  • Grant proposal (Peter/Martin) - Gordon / Lori - October 2015
    • 100714
    • proposal submitted. Now waiting for results.

Ripple

  • Hardware upgrade - reconfigure RAID (Peep) - Lori
    • completed. now more than twice as fast
  • CRS - reconfigured, adding "clucks" + 40Gb switch - Lori
    • installed, re-organized into two clusters. Will configure SALT to manage systems
  • 100235

SciCom

  • Upgrade elora.cs to 14.04 - Mike
    • need ST item

SWAG

  • Upgrade swag.cs to 14.04 - Mike

Watform

  • Retire old servers - Ronaldo
    • done
  • Documentation - Ronaldo
Topic revision: r12 - 2015-12-16 - LawrenceFolland
 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback